OpenStudy (anonymous):

suppose y varies directly with x write a direct variation equation that relates x and y. y=15 when x=6

OpenStudy (anonymous):

according to question, y = k x, where k= constant now, 15 = k 6 k = 5/2 so, y = 5/2 x 2y =5x